# The n_TOF Collaboration

### v1

@@ Abstract
Background
50,53Cr are very relevant in criticality safety benchmarks related to nuclear reactors. The discrepancies of up to 30% between the neutron capture cross section evaluations have an important effect on the $k_{eff}$ and $k_{\infty}$ in criticality benchmarks particularly sensitive to chromium.
Purpose
The 50,53Cr(n,$\gamma$) cross sections is to be determined between 1 and 100 keV with an 8-10% accuracy following the requirements of the NEA High Priority Request List (HPRL) to solve the current discrepancies.
Method
We have measured the neutron capture cross sections by the time-of-flight technique at the EAR1 experimental area of the n_TOF facility, using an array of four C6D6 detectors with very low neutron sensitivity. The highly-enriched samples used are significantly thinner than in previous measurements, thus minimizing the multiple-scattering effects.
Results
We have produced, and analyzed with the R-matrix analysis code SAMMY, capture yields featuring 33 resonances of 50Cr and 51 of 53Cr with an accuracy between 5% and 9%, hence fulfilling the requirements made by the NEA. The differential and integral cross sections have been compared to previous data and evaluations.
Conclusions
The new 50,53Cr(n,$\gamma$) cross sections measured at the CERN n_TOF facility provide a valuable input for upcoming evaluations, which are deemed necessary given that the results presented herein do not support the increase in both cross sections proposed in the recent INDEN evaluation.
preprint: APS/123-QED
I Introduction
@@
In nuclear reactors, chromium is an important component (11-26% abundance) of the stainless steel used as a structural material, and thus its neutron-interaction cross section plays an important role in the associated neutronic calculations. In particular, Trkov [3] reported significant discrepancies between calculations and integral measurements when performing $k_{eff}$ and $k_{\infty}$ criticality benchmarks sensitive to chromium like HEU-COMP-INTER-005/4=KBR-15/Cr or PU-MET-INTER-002. These are due to the discrepancies of about 30% between the chromium cross sections libraries/evaluations. These findings led to a new entry in the Nuclear Energy Agency (NEA) High Priority Request List (HPRL) [4] calling for a new neutron capture cross section measurement of 50Cr and 53Cr between 1 and 100 keV with an accuracy of 8-10%.
@@
The need of new measurements is related to the limited accuracy and experimental issues of the previous ones. Chromium is a relatively light isotope, and it is characterized by its large scattering-to-capture cross section ratio, especially at neutron energies $E_{n}$ between 1 and 10 keV, where several resonances with a large scattering cross section are found. They are sometimes referred to as a s-wave resonance cluster. Therefore, the neutron capture and transmission measurements of 50,53Cr made with relatively thick samples in the last decades of the XX century  [5, 6, 7, 8] suffered from high sensitivity to scattered neutrons and strong multiple-scattering effects. More recently, in 2011 a neutron capture and transmission measurement of natural and 53Cr was performed by Guber et al. [9] at ORNL ORELA using low neutron sensitivity detectors, but the multiple-scattering was still an issue, as discussed later by Nobre et al. [10]. Considering the capture and transmission data from Ref. [9] together with another (unpublished) ORELA transmission data from Harvey et al., Leal et al. [11] produced a cross section evaluation that has been adopted in JEFF-3.3 [12], ENDF/B-VIII.0 [13], JENDL-5 [14] and BROND-3.1 [15]; while CENDL-3.2 [16] is based on the data published before 2011. Only a few years ago the new INDEN evaluation [17] has been released (and adopted in ENDF/B-VIII.1 [18]), proposing a significant increase of the capture cross section of both 50Cr and 53Cr based on a re-analysis by Nobre et al. [10] of the data from Stieglitz et al. [5] and Guber et al. [9] in which issues with the sample thickness are identified and addressed through detailed MCNP based multiple-scattering corrections. This new evaluation provide a better performance on aforementioned criticality benchmarks. Overall, the discrepancies in the chromium cross sections in the evaluations have a large impact of about 1000 pcm (or a 1%) in the mentioned criticality calculations.
@@
In this context, and in response to the NEA HPRL request, two measurements have been designed and performed: a time-of-flight measurement at n_TOF-EAR1 carried out during summer 2022, and a 50Cr neutron activation measurement at the HiSPANoS facility of CNA [19, 20] performed in winter 2023, in which the Maxwellian Averaged Cross Section at $kT=30$ keV $\text{MACS}_{30}$ of 50Cr has been determined experimentally for the first time [21]. In this work we describe and discuss the n_TOF measurement, of which preliminary results were presented in the WONDER 2023 Workshop (see Ref. [22]). The experimental set-up is described in Sec. II, the extraction of the capture yield is detailed in Sec. III, and the R-matrix resonance analysis and the results obtained are described in Sec. IV. Last, the conclusions of this work are summarized in Sec. V.
II Measurement at n_TOF
II.1 The n_TOF facility at CERN
@@
II.2 Chromium and ancillary samples
@@
As neutron multiple-scattering effects are believed to be the main issues in previous experiments, the samples for this measurement were designed to be much thinner that those in all previous cases, thus minimizing these effects. Two samples were made for each isotope: a very thin one to measure between 1 and 10 keV, and a thicker one, but still thinner than those used in previous experiments, to cover the range from 10 to 100 keV with enough statistics. The physical characteristics of each sample are summarized in Table 2, where the mass reported corresponds to the total amount of chromium oxide, from which the atomic thickness $n_{\text{at}}$ corresponding to the chromium atoms is calculated assuming the isotopic composition and uncertainty provided by the supplier of the material.
The powder was pressed into a pellet inside a 20 mm diameter, 0.5 mm thick and 4.5 mm tall (or 11.5 mm for the 53Cr-thick sample) PEEK capsule with the top closed by a thin Mylar layer. This design minimizes neutron capture in the capsule, allows for a regular visual inspection of the pressed pellet to verify its integrity (as it needed to be moved in and out of the beam several times during the irradiation), and avoids losing material if the pellet would break. A series of X-ray Computed Tomographies (CT) were performed by the MME Group at CERN to each chromium sample to measure their thickness (expressed in mm in Table 2) and identify any significant inhomogeneity. The CT of the 50Cr-thick sample is shown in Fig. 1 as an example, where local variations of the thickness and a slightly concave meniscus at its edges can be observed. Monte Carlo simulations have shown that these aspects did not affect significantly the extraction of the capture cross section [22]. The samples were placed in the center of a 50 mm diameter aluminium ring, glued on a Mylar foil, which is the conventional sample holder for neutron capture experiments at n_TOF EAR1. The 53Cr-thick sample mounted in the measuring position is shown in Fig. 2.
@@
II.3 Experimental set-up
The neutron capture detection set-up consisted of an array of four C6D6 detectors [28] with a cylindrical active volume of 1 L, encapsulated in a carbon fiber housing. This design by Plag et al. [29] reduces the neutron sensitivity of the detectors, which is needed to reduce as much as possible the background related to the large neutron scattering cross section in chromium. They also feature a very small gamma-ray detection efficiency, an attribute needed to apply the Pulse Height Weighting Technique (PHWT) [30] to asses the efficiency for detecting capture cascades (see Sec. III.3). The detectors were placed at 8 cm from the sample position, at a 125 degrees backward angle with respect to the neutron beam, as shown in Fig. 2, to reduce the background produced by in-beam photons undergoing Compton scattering in the sample, and to minimize possible anisotropies in the gamma-ray emission for resonances with orbital angular momentum $\ell>0$.
@@ II.3 Experimental set-up
III Determination of the capture yield
@@
where $C_{w}$ and $B_{w}$ are the total and background weighted counting rates (neglecting dead-time corrections because of the low counting rates), $E_{c}$ the capture cascade energy (see Sec. III.3), $\Phi$ the neutron flux at EAR1, $F_{BIF}$ the fraction of the neutron flux seen by the sample, and $F_{PHWT}$ a correction factor associated with the PHWT. The analysis to obtain all these quantities is detailed in the following sections.
III.1 Detector and time-of-flight calibration
The detectors were calibrated once per week both in $\gamma$-ray deposited energy and resolution using radioactive sources of 137Cs, 88Y and 241Am-9Be which emit $\gamma$-rays of 0.662, 0.898 and 1.836, and 4.438 MeV, respectively. The end point of 197Au capture cascade at 6.512 MeV was also included in order to extend the deposited energy range. These measurements were complemented with GEANT4 simulations [37, 38] of the C6D6 detectors response, in which a realistic model of the experimental area and the whole capture set-up was implemented. A small gain shift ($\sim 2\%~{}$ over the full campaign) was observed, especially on one of the detectors, and thus the experiment was divided into six periods, each with its own energy calibration.
@@ III.1 Detector and time-of-flight calibration
The neutron energy was calibrated using the capture yield of 197Au, whose resonance energies are well known below 2 keV. The yield was analyzed with the R-matrix Bayesian code SAMMY [42] (see Sec. IV), including a numerical version of the n_TOF RF obtained from simulations. The nominal resonance energies found in JEFF-3.3 were represented against the reconstructed values from the time-of-flight, as shown in Fig. 3. Then, a simultaneous fit of $L$ and $t_{off}$ was performed, obtaining $L=183.94(4)~{}$m and $t_{off}=-17.57(4)~{}$ns, which allowed to reproduce the energy of the 197Au resonances within 0.1%.
@@
III.2 Background determination
As mentioned in Sec. II.2, a series of measurements with auxiliary samples were performed in order to estimate the background. The different sources of background can be classified depending on whether their origin is sample-related or sample-independent.
@@ III.2 Background determination
The total background depicted in Fig. 4 is thus considered a good estimation and hence was subtracted. The remaining smooth background component is included in the resonance analysis with SAMMY [42] (see Sec. IV).
@@
III.3 Total Energy Detection technique
@@
These effects have been considered simultaneously [44], estimating $F_{PHWT}$ by simulating the response of the detectors not to individual $\gamma$-rays but to capture cascades emitted from each sample. This requires a set of realistic capture cascades, which have been obtained with the cascade generator software NuDEX [47]. NuDEX takes the level scheme and the branching ratios below a certain excitation energy from ENSDF [48]. At higher energies, where the levels and branching ratios are not known, they are randomly generated according to statistical models of the nuclear level density. Each set of levels and branching ratios between all of them but the initial state is called a realization. In practice, the de-excitation path starts at a specific resonance, and the branching of individual resonances differs. The specific choice of this branching for a fixed realization is called a subrealization.
@@
Table 3: Correction factors $F_{PHWT}$ for each individual sample and detector. The systematic uncertainty of the individual values is 1.7% (see text for details, is not included).
@@
III.4 Normalization through the Saturated Resonance Method
The Beam Interception Factor (BIF) is defined as the fraction of neutrons in the beam seen by the sample, which is introduced in the calculation of the capture yield as a normalization factor $F_{BIF}$. This factor is obtained through the so-called Saturated Resonance Method (SRM) [27] using the 4.9 eV resonance of 197Au measured with a thick target and thus featuring a saturation of the yield, demonstrated as a flat resonance top, as all neutrons impinging the sample are captured (see Fig. 6). The capture yield of the 20 mm 197Au sample has been obtained for each detector and analysed with SAMMY [42] leaving the normalization factor free to vary. The resulting normalization factors provide the values of $F_{BIF}$, which agree within 1.4% for individual detectors, yielding average values of $F_{BIF}=0.645(9)$. The same procedure was performed to the capture yield of the 80 mm diameter gold sample, which is larger than the beam diameter, obtaining 1.003(3), in a perfect agreement with the expected value of 1.
@@ III.4 Normalization through the Saturated Resonance Method
III.5 Capture yield: PWHT implementation for 50,53Cr
The implementation of the PHWT enhances the statistical fluctuations of the data because the weights are assigned in a signal-by-signal basis, something that has been studied in detail recently by Mendoza et al. [49]. Having limited statistics when applying the PHWT is equivalent to sampling poorly the detector response. This becomes critical when the statistics are very limited, causing the overall weighting to vary significantly between neighbouring bins. As a consequence, fluctuations on individual points are enhance, inducing deformations in the shape of the resonances after the weighting process. In our case, this issue is severely affecting the yield above $\sim$20 keV due to the limited statistics in combination with a significant contribution of high-energy $\gamma$-rays to chromium spectra.
As shown with detail in Ref [50], we can suppress the fluctuations in weighted counts $C_{w}$ by defining energy regions where the detection efficiency is constant. We have developed a new method in which, instead of applying directly the PHWT calculated in narrow neutron energy bins, we determine a weighting factor $WF_{res}$ independently for each resonance: i.e. a resonance weighting factor (RWF) is applied. The $WF_{res}$ factors are obtained by computing the ratio between the area of each resonance in the unweighted and weighted yields. This method is a generalization of the Average Weighting Factor (AWF) technique introduced in Ref. [51]. Our method is useful for nuclei where the detection efficiency changes significantly between individual resonances, as in the case of chromium, so it works very well for isolated resonances. Obviously, the RWF can only be calculated for resonances with enough statistics so $WF_{res}$ can be accurately extracted, and otherwise the AWF had to be used instead, as it is done in the valleys between resonances. We have included an additional systematic uncertainty in the capture yield, corresponding to the statistics of each resonance as it affects the $WF_{res}$. This is shown in Fig. 7. Since the statistics of the main s-wave resonances located between 1 and 10 keV are much higher, they are not significantly affected by the fluctuations, so the yield in that region has been obtained applying the "standard" PHWT, and thus is not affected by this additional uncertainty.
@@ III.5 Capture yield: PWHT implementation for 50,53Cr
  • For the range between 10 and 100 keV, the capture yield has been extracted from the thick samples. Parameters of 42 resonances have been analyzed using the new RWF technique, and the remaining ones (with poor statistics) along with the data between resonances have been analyzed by the AWF technique. The capture yield in this region has an additional systematic uncertainty because of this procedure (see Table 4).
III.6 Systematic uncertainty
The different sources of systematic uncertainties have been described throughout the text, and are summarized in Table 4. The estimation of uncertainties is conservative but still yields an overall accuracy between 5% and 9% in the energy range of interest, between 1 and 100  keV. This is indeed within the 8-10% requested in the NEA High Priority Request List (HPRL) motivating this experiment.
Table 4: Contributions to the systematic uncertainty of the capture yields. As described in Sec. III.5, the RWF correction applies only to neutron energies above 10 keV.

### v2

@@ Abstract
[…]
1 Introduction
@@
In nuclear reactors, chromium is an important component (11-26% abundance) of the stainless steel used as a structural material, and thus its neutron-interaction cross section plays an important role in the associated neutronic calculations. In particular, Trkov [3] reported significant discrepancies between calculations and integral measurements when performing $k_{eff}$ and $k_{\infty}$ criticality benchmarks sensitive to chromium like HEU-COMP-INTER-005/4=KBR-15/Cr or PU-MET-INTER-002. These are due to discrepancies of about 30% between the chromium cross sections libraries/evaluations found at the time. These findings led to a new entry in the Nuclear Energy Agency (NEA) High Priority Request List (HPRL) [4] calling for a new neutron capture cross section measurement of 50Cr and 53Cr between 1 and 100 keV with an accuracy of 8-10%.
@@
The need of new measurements is related to the limited accuracy and experimental issues of the previous ones. Chromium is a relatively light isotope, and it is characterized by its large scattering-to-capture cross section ratio, especially at neutron energies $E_{n}$ between 1 and 10 keV, where several resonances with a large scattering cross section are found. They are sometimes referred to as a s-wave resonance cluster. Therefore, the neutron capture and transmission measurements of 50,53Cr made with relatively thick samples in the last decades of the XX century  [5, 6, 7, 8] suffered from high sensitivity to scattered neutrons and strong multiple-scattering effects. More recently, in 2011 a neutron capture and transmission measurement of natural and 53Cr was performed by Guber et al. [9] at ORNL ORELA using low neutron sensitivity detectors, but the multiple-scattering was still an issue, as discussed later by Nobre et al. [10]. Considering the capture and transmission data from Ref. [9] together with another (unpublished) ORELA transmission data from Harvey et al., Leal et al. [11] produced a cross section evaluation that has been adopted in JEFF-3.3 [12], ENDF/B-VIII.0 [13], JENDL-5 [14] and BROND-3.1 [15]; while CENDL-3.2 [16] is based on the data published before 2011. Only a few years ago the new INDEN evaluation [17] has been released (and adopted in ENDF/B-VIII.1 [18]), proposing a significant increase of the capture cross section of both 50Cr and 53Cr based on a re-analysis by Nobre et al. [10] of the data from Stieglitz et al. [5] and Guber et al. [9] in which issues with the sample thickness are identified and addressed through detailed MCNP based multiple-scattering corrections. This new evaluation provide a better performance on aforementioned criticality benchmarks. Overall, the discrepancies in the chromium cross sections in the evaluations have a large impact of about 1000 pcm (or 1%) in the mentioned criticality calculations.
@@
In this context, and in response to the NEA HPRL request, two measurements have been designed and performed: a time-of-flight measurement at n_TOF-EAR1 carried out during summer 2022, and a 50Cr neutron activation measurement at the HiSPANoS facility of CNA [19, 20] performed in winter 2023, in which the Maxwellian Averaged Cross Section at $kT=30$ keV ($\text{MACS}_{30}$) of 50Cr has been determined experimentally for the first time [21]. In this work we describe and discuss the n_TOF measurement, of which preliminary results were presented in the WONDER 2023 Workshop (see Ref. [22]). The experimental set-up is described in Sec. 2, the extraction of the capture yield is detailed in Sec. 3, and the R-matrix resonance analysis and the results obtained are described in Sec. 4. Last, the conclusions of this work are summarized in Sec. 5.
2 Measurement at n_TOF
2.1 The n_TOF facility at CERN
@@
2.2 Chromium and ancillary samples
@@
As neutron multiple-scattering effects are believed to be the main issues in previous experiments, the samples for this measurement were designed to be much thinner that those in all previous cases, thus minimizing these effects. Two samples were made for each isotope: a very thin one to measure between 1 and 10 keV, and a thicker one, but still thinner than those used in previous experiments, to cover the range from 10 to 100 keV with enough statistics. The physical characteristics of each sample are summarized in Table 2, where the mass reported corresponds to the total amount of chromium oxide, from which the atomic thickness $n_{\text{at}}$ equivalent to the chromium atoms is calculated assuming the isotopic composition and uncertainty provided by the supplier of the material.
The powder was pressed into a pellet inside a 20 mm diameter, 0.5 mm thick and 4.5 mm tall (or 11.5 mm for the 53Cr-thick sample) PEEK capsule with the top closed by a thin Mylar layer. This design minimizes neutron capture in the capsule, allows for a regular visual inspection of the pressed pellet to verify its integrity (as it needed to be moved in and out of the beam several times during the irradiation), and avoids losing material if the pellet would break. A series of X-ray Computed Tomographies (CT) were performed by the MME Group at CERN to each chromium sample to measure their thickness (expressed in mm in Table 2) and identify any significant inhomogeneity. The CT of the 50Cr-thick sample is shown in Fig. 1 as an example, where local variations of the thickness and a slightly concave meniscus at its edges can be observed. Monte Carlo simulations have shown that these aspects did not affect significantly the extraction of the capture cross section [22]. The samples were placed in the centre of a 50 mm diameter aluminium ring, glued on a Mylar foil, which is the conventional sample holder for neutron capture experiments at n_TOF EAR1. The 53Cr-thick sample mounted in the measuring position is shown in Fig. 2.
@@
2.3 Experimental set-up
The neutron capture detection set-up consisted of an array of four C6D6 detectors [28] with a cylindrical active volume of 1 L, encapsulated in a carbon fibre housing. This design by Plag et al. [29] reduces the neutron sensitivity of the detectors, which is needed to reduce as much as possible the background related to the large neutron scattering cross section in chromium. They also feature a very small gamma-ray detection efficiency, an attribute needed to apply the Pulse Height Weighting Technique (PHWT) [30] to asses the efficiency for detecting capture cascades (see Sec. 3.3). The detectors were placed at 8 cm from the sample position, at a 125 degrees backward angle with respect to the neutron beam, as shown in Fig. 2, to reduce the background produced by in-beam photons undergoing Compton scattering in the sample, and to minimize possible anisotropies in the gamma-ray emission for resonances with orbital angular momentum $\ell>0$.
@@ 2.3 Experimental set-up
3 Determination of the capture yield
@@
where $C_{w}$ and $B_{w}$ are the total and background weighted counting rates (neglecting dead-time corrections because of the low counting rates), $E_{c}$ the capture cascade energy (see Sec. 3.3), $\Phi$ the neutron flux at EAR1, $F_{BIF}$ the fraction of the neutron flux seen by the sample, and $F_{PHWT}$ a correction factor associated with the PHWT. The analysis to obtain all these quantities is detailed in the following sections.
3.1 Detector and time-of-flight calibration
[…]
@@ 3.1 Detector and time-of-flight calibration
The neutron energy was calibrated using the capture yield of 197Au, whose resonance energies are well known below 2 keV. The yield was analysed with the R-matrix Bayesian code SAMMY [42] (see Sec. 4), including a numerical version of the n_TOF RF obtained from simulations. The nominal resonance energies found in JEFF-3.3 were represented against the reconstructed values from the time-of-flight, as shown in Fig. 3. Then, a simultaneous fit of $L$ and $t_{off}$ was performed, obtaining $L=183.94(4)~$m and $t_{off}=-17.57(4)~$ns, which allowed to reproduce the energy of the 197Au resonances within 0.1%.
@@
3.2 Background determination
As mentioned in Sec. 2.2, a series of measurements with auxiliary samples were performed in order to estimate the background. The different sources of background can be classified depending on whether their origin is sample-related or sample-independent.
@@ 3.2 Background determination
The total background depicted in Fig. 4 is thus considered a good estimation and hence was subtracted. The remaining smooth background component is included in the resonance analysis with SAMMY [42] (see Sec. 4).
@@
3.3 Total Energy Detection technique
@@
These effects have been considered simultaneously [44], estimating $F_{PHWT}$ by simulating the response of the detectors not to individual $\gamma$-rays but to capture cascades emitted from each sample. This requires a set of realistic capture cascades, which have been obtained with the cascade generator software NuDEX [47]. NuDEX takes the level scheme, Photon Strength Functions and the branching ratios below a certain excitation energy from ENSDF [48] and RIPL-3 [49]. At higher energies, where the levels and branching ratios are not known, they are randomly generated according to statistical models of the nuclear level density. Each set of levels and branching ratios between all of them but the initial state is called a realization. In practice, the de-excitation path starts at a specific resonance, and the branching of individual resonances differs. The specific choice of this branching for a fixed realization is called a subrealization.
@@
Table 3: Correction factors $F_{PHWT}$ for each individual sample and detector. The systematic uncertainty of the individual values is 1.7% (see text for details).
@@
3.4 Normalization through the Saturated Resonance Method
The Beam Interception Factor (BIF) is defined as the fraction of neutrons in the beam seen by the sample, which is introduced in the calculation of the capture yield as a normalization factor $F_{BIF}$. This factor is obtained through the so-called Saturated Resonance Method (SRM) [27] using the 4.9 eV resonance of 197Au measured with a thick target and thus featuring a saturation of the yield, demonstrated as a flat resonance top, as all neutrons impinging the sample are captured (see Fig. 6). The capture yield of the 20 mm 197Au sample has been obtained for each detector and analysed with SAMMY [42] leaving the normalization factor free to vary. The resulting normalization factors provide the values of $F_{BIF}$, which agree within 1.4% for individual detectors, yielding an average value of $F_{BIF}=0.645(9)$. The same procedure was performed to the capture yield of the 80 mm diameter gold sample, which is larger than the beam diameter, obtaining 1.003(3), in a perfect agreement with the expected value of 1.
@@ 3.4 Normalization through the Saturated Resonance Method
3.5 Capture yield: PWHT implementation for 50,53Cr
The implementation of the PHWT enhances the statistical fluctuations of the data because the weights are assigned in a signal-by-signal basis, something that has been studied in detail recently by Mendoza et al. [50]. Having limited statistics when applying the PHWT is equivalent to sampling poorly the detector response. This becomes critical when the statistics are very limited, causing the overall weighting to vary significantly between neighbouring bins. As a consequence, fluctuations on individual points are enhanced, inducing deformations in the shape of the resonances after the weighting process. In our case, this issue is severely affecting the yield above $\sim$20 keV due to the limited statistics in combination with a significant contribution of high-energy $\gamma$-rays to chromium spectra.
As shown with detail in Ref. [51], we can suppress the fluctuations in weighted counts $C_{w}$ by defining energy regions where the detection efficiency is constant. We have developed a new method in which, instead of applying directly the PHWT calculated in narrow neutron energy bins, we determine a weighting factor $WF_{res}$ independently for each resonance: i.e. a resonance weighting factor (RWF) is applied. The $WF_{res}$ factors are obtained by computing the ratio between the area of each resonance in the unweighted and weighted yields. This method is a generalization of the Average Weighting Factor (AWF) technique introduced in Ref. [52]. Our method is useful for nuclei where the detection efficiency changes significantly between individual resonances, as in the case of chromium, so it works very well for isolated resonances. Obviously, the RWF can only be calculated for resonances with enough statistics so $WF_{res}$ can be accurately extracted, and otherwise the AWF had to be used instead, as it is done in the valleys between resonances. We have included an additional systematic uncertainty in the capture yield, corresponding to the statistics of each resonance as it affects the $WF_{res}$. This is shown in Fig. 7. Since the statistics of the main s-wave resonances located between 1 and 10 keV are much higher, they are not significantly affected by the fluctuations, so the yield in that region has been obtained applying the "standard" PHWT, and thus is not affected by this additional uncertainty.
@@ 3.5 Capture yield: PWHT implementation for 50,53Cr
  • For the range between 10 and 100 keV, the capture yield has been extracted from the thick samples. Parameters of 42 resonances have been analysed using the new RWF technique, and the remaining ones (with poor statistics) along with the data between resonances have been analysed by the AWF technique. The capture yield in this region has an additional systematic uncertainty because of this procedure (see Table 4).
3.6 Systematic uncertainty
The different sources of systematic uncertainties have been described throughout the text, and are summarized in Table 4. The estimation of uncertainties is conservative but still yields an overall accuracy between 5% and 9% in the energy range of interest (1-100 keV). This is indeed within the 8-10% requested in the NEA High Priority Request List (HPRL) motivating this experiment.
Table 4: Contributions to the systematic uncertainty of the capture yields. As described in Sec. 3.5, the RWF correction applies only to neutron energies above 10 keV.